# Gravitational waves from a dilaton-induced, first-order QCD phase transition

### v1

@@ Abstract
where $\alpha_{s}(\phi)=1/(4\pi\mathfrak{f}(\phi))$ 222 This correction is valid for the $\phi$ values we are interested in at temperatures higher than the confinement scale, since the RG-evolved coupling satisfies $\alpha_{s}(\phi_{\rm TV},\mu=T)<1$, so that QCD is still perturbative.. We note that the interactions of dilaton fluctuations with the gluons with vertex structure $\sim(p_{1}^{\mu}p_{2}^{\nu}-g^{\mu\nu})$ renormalise the couplings but do not have finite contributions at $T\neq 0$ that could change the shape of the 1-loop thermal potential, and hence we neglect these contributions.
The effective potential of Eq. (6) is shown schematically in the left panel of Fig. 1 for four different temperatures. At sufficiently high temperatures ($T\gg m_{\phi}$), $V_{\rm eff}$ is dominated by the stabilising, leading-order contribution from $V_{T}$, $\sim m^{2}(\phi)T^{2}$, and the loop-corrected part of $V_{\cal P}$, which push the dilaton towards weak coupling (corresponding to larger field values). This leads to a single minimum located at $\phi_{\rm TV}<\phi<\phi_{\rm FV}$.
At intermediate temperatures, the thermal contributions from self-interactions become Boltzmann-suppressed and can be ignored, and $V_{\rm eff}$ develops a second minimum at smaller field values. Notably, due to $V_{\cal P}$, $\phi_{\rm FV}$ is the global minimum of $V_{\rm eff}$ at these temperatures, and the dilaton is attracted to the false vacuum.
@@
The subsequent cosmology and observational predictions of this scenario depend on the speed at which the bubbles expand, which we now discuss.
Bubble wall velocity — We require that the transition percolates (at plasma temperature $T_{\star}$), which for a PT with constant $\Gamma$ during the hot Big Bang cosmology requires $\alpha\lesssim 20$ [33], where $\alpha=\Delta V/\rho_{\rm rad}(T_{\rm n})$. For transitions with time-independent $\Gamma$, the corresponding $\beta$-parameter that sets the inverse timescale of the PT is given by $\beta=\dot{I}(t_{\star})$, with $I(t)$ the average number of bubbles in the past comoving light cone [33], i.e.
@@
where $a(t)$ denotes the scale factor, $r_{\rm com}(t_{\star},t)$ the comoving bubble radius and $t_{\star}$ the percolation time, defined by $I(t_{\star})=1$. For $0<\alpha\lesssim 20$, this gives $38\gtrsim\beta>$ 3, correspondingly. As the energy density $\Delta V$ is channelled into the Standard Model, the plasma is reheated to the temperature $T_{\rm rh}=(1+\alpha)^{1/4}T_{n}$, here assuming $T_{\star}\approx T_{n}$. We restrict our discussion to $T_{\rm rh}\lesssim T_{\rm QCD}^{\rm TV}$, so that the hadrons do not deconfine (and then undergo a ‘second’ QCD transition, this time a cross-over).
The bubble wall velocity is determined by the balance of the driving pressure, $\mathcal{P}_{\rm dr}\equiv\Delta V$, and the pressure caused by friction from the surrounding plasma, $\mathcal{P}_{\rm fr}$. The calculation of bubble wall velocities in confining, first-order PTs is an active area of research [34, 35, 36, 36, 37, 38]. In this work, the driving pressure is generated by the dilaton potential instead of the confinement dynamics. The expanding bubbles are in the ballistic limit, meaning that the particles in the plasma do not interact significantly over the width of the wall, which in the plasma frame is given by $L_{w}/\gamma_{w}\sim 1/(m_{\phi}\gamma_{w})$ where $\gamma_{w}$ the Lorentz factor of the wall. The ballistic limit holds when $L_{w}/\gamma_{w}\ll L_{\rm MFP}$, where $L_{\rm MFP}$ is the mean-free path of the particles in the quark-gluon plasma outside the bubbles [39]. Parametrically, we expect $L^{-1}_{\rm MFP}\sim\alpha^{2}_{\rm s}(\phi_{\rm FV})\;T_{n}$ from $2\rightarrow 2$ scatterings. Since QCD remains perturbative in the false vacuum, the ballistic approximation holds for all values of $\gamma_{w}$.
$\mathcal{P}_{1\to 1}=\frac{T_{n}^{2}}{48}\left(g_{q}^{*}\Delta m_{q}^{2}+2g_{g}^{*}\Delta m_{g}^{2}\right)<\alpha_{N}\rho_{\rm rad}=\mathcal{P}_{\rm dr}\;,$ (10)
where $g_{q}^{*},\,g^{*}_{g}$ denote, respectively, the effective degrees of freedom for quarks and gluons in the plasma. The inequality assumes $\alpha\gtrsim 3\cdot 10^{-3}\,(\Lambda_{\rm QCD}/T_{n})^{2}$, which holds across the most relevant parameter space. Thus, one-to-one processes alone do not suffice to stop the runaway behaviour [40] 333 A more accurate description of friction in a confinement phase transition is given in [78], and results in the scaling $\mathcal{P}_{\rm friction}=\mathcal{P}_{\rm BS}\sim\sum_{i\in BS}\frac{m_{i}^{2}T_{n}^{2}}{24}$, which coincides with that of [40], used here. .
Gluon radiation is an additional source of friction. Here, we model the pressure from gluon emission at the wall interface by the equivalent expression for $W$ and $Z$ boson emission when entering bubbles of the low-temperature phase in models where the electroweak PT is first order [42] (we explore an alternative scenario with runaway walls in the Appendix). Adopting the results of [42, 43], the gluon emission pressure is given by $\mathcal{P}_{1\rightarrow 2}\simeq N_{g}\Delta m_{g}\gamma\;\alpha_{s}(\phi_{w})T_{n}^{3}$, with $\phi_{w}$ a value of the dilaton on the wall and $N_{g}=8$ is the number of gluons. Here, we have assumed that QCD remains perturbative at the wall, i.e. for an intermediate value of $\phi$. This is typically justified as $\Lambda_{\rm QCD}(\phi_{w})\ll\Lambda_{\rm QCD}$, by Eq. (3). The pressure contribution $\mathcal{P}_{1\rightarrow 2}$ grows with $\gamma$ and eventually catches up with $\Delta V$, leading to an ultra-relativistic terminal velocity with gamma factor  [^4] Another potential friction source is bound state formation from string fragmentation[34]. However, due to insignificant supercooling, we expect confinement to happen among quark pairs rather than quarks forming flux-tubes attached to the bubble wall, and string fragmentation to be unimportant. Even in this scenario though, the leading and next to leading order friction from gluons has similar dependence on $\gamma$
By comparing the size of the bubbles when reaching the terminal velocity to their separation, it is clear that the terminal velocity is reached well before bubbles collide.
Gravitational wave signal — As the relativistic dilaton bubbles grow large, collide, and percolate, they produce a stochastic gravitational wave (GW) background, potentially observable today [45, 46]. This is particularly interesting given the recent observation of the Hellings-Down correlation in Pulsar Timing Array (PTA) datasets [4, 5, 6, 7], providing evidence for a stochastic GW background in the nanohertz range. Such a background could potentially be explained astrophysically through supermassive black hole binary mergers [47], but this explanation has difficulties with accommodating the observed amplitude and spectrum [48, 49]. By contrast, several studies have found that GWs generated by a first-order PT provides a better fit to the data [50, 51, 52, 53, 54, 55, 56] (see [57] for an alternative gravitational wave explanation generated by turbulent sources regardless of the order of the QCD PT).
The main sources of the GW background in our model are the expansion and collision of bubble walls [58, 59, 60], and the dynamics of sound waves generated in the plasma [61, 62, 63]. In both cases, estimates of the amplitude and spectrum of the GW background can be parametrised through rather simple formulas, depending on $T_{\star}\approx T_{n},~{}\alpha,~{}\beta$, and $v_{w}$ (cf. [43] and, for a recent review, [64]). The peak frequency is given by
$f_{i}=c_{i}(v_{w})\frac{\beta}{H}\frac{T_{n}}{100~{}\text{MeV}}\left(\frac{g_{*}}{61.75}\right)^{1/6}\;,$ (12)
where $i\in\{{\rm coll.},~{}{\rm sw}\}$ for the collision and sound wave contributions, respectively. The velocity-dependent prefactors are given by [65]
$c_{\rm coll.}=\frac{9.44~{}\text{nHz}}{1.8-0.1v_{w}+v_{w}^{2}}\;,~{}c_{\rm sw}=\frac{17.53~{}\text{nHz}}{v_{w}}\,.$ (13)
[…]
@@
The corresponding GW energy spectra, $h^{2}\Omega_{\rm GW}$, give the fractional energy density per logarithmic frequency interval, rescaled by the squared Hubble parameter. Since the bubbles are expected to reach a terminal velocity before they collide, the main contribution to the GW signal in this scenario is from sound waves [65, 66] (the contribution to the signal coming from bubble wall collisions, which is relevant for runaway PTs, is discussed in the Appendix). The GW spectrum from sound waves is then given by
@@
Acknowledgments.–We thank Mark Hindmarsh and Filippo Sala for stimulating discussions. This work was supported by the Swedish Research Council (VR) under grants 2018-03641, 2019-02337 and 2024-04289. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
Appendix A Appendix: Runaway walls
@@
In the bulk of this paper, we have discussed the well-motivated scenario where the dilaton bubble walls reach a terminal velocity before colliding. It is instructive to explore the alternate scenario in which the bubble walls ‘run away’. Such a scenario can be realised in models that yield sufficiently small QCD coupling at the bubble wall interface along with large values of $\beta/H$ so that bubbles never reach their terminal velocity before collision. When this happens, the dominant contribution to the gravitational wave spectrum is given by bubble wall collisions. As discussed, quantum phase transitions with $\Gamma\simeq$ constant yield $\beta/H\lesssim 8$, which is not very large. However, if the dilaton phase transition is for some reason driven by thermal fluctuations, larger values of $\beta/H$ can be attained, triggering runaway behaviour.
@@
Figure 3: Same as Fig. 2 but including the contribution from bubble wall collisions. Solid lines show the spectra in the regime where the energy is carried by bubble walls, while dashed lines correspond to the regime where the energy is converted into sound waves, for comparison.
@@
$h^{2}\Omega_{\rm coll.}(f)=1.96\times 10^{-5}\left(\frac{H}{\beta}\right)^{2}\left(\frac{\kappa_{\rm coll.}\alpha}{1+\alpha}\right)^{2}$
×$\times\left(\frac{61.75}{g_{*}(T_{\star})}\right)^{1/3}\left(\frac{0.11v_{w}^{3}}{0.42+v_{w}^{2}}\right)\times\frac{3.8\left(f/f_{\rm coll.}\right)^{2.8}}{1+2.8\left(f/f_{\rm coll.}\right)^{3.8}}\;.$
Here $v_{w}=1$ and $\kappa_{\rm coll}$ is energy fraction that is converted into bubble walls. The corresponding spectra are shown in Fig. 3, where we use the same parameter values as in Fig. 2, but now assume that all of the energy is in the form of colliding bubble walls, i.e. $\kappa_{\rm coll}=1$. The dashed lines in Fig. 3 show the corresponding spectra from Fig. 2, arising from sound waves.

### v2

@@ Abstract
where $\alpha_{s}(\phi)=1/(4\pi\mathfrak{f}(\phi))$ 222 This correction is valid for the $\phi$ values we are interested in at temperatures higher than the confinement scale, since the RG-evolved coupling satisfies $\alpha_{s}(\phi_{\rm TV},\mu=T)<1$, so that QCD is still perturbative.,  [^3] We note that the interactions of dilaton fluctuations with the gluons with vertex structure $\sim(p_{1}^{\mu}p_{2}^{\nu}-g^{\mu\nu})$ renormalise the couplings but do not have finite contributions at $T\neq 0$ that could change the shape of the 1-loop thermal potential, and hence we neglect these contributions..
The effective potential of Eq. (6) is shown schematically in the left panel of Fig. 1 for four different temperatures. At sufficiently high temperatures ($T\gg m_{\phi}$), $V_{\rm eff}$ is dominated by the stabilising, leading-order contribution from $V_{T}$, $\sim m^{2}(\phi)T^{2}$, and the loop-corrected part of $V_{\cal P}$, which push the dilaton towards weak coupling (corresponding to larger field values). The potential has a single minimum in the range $\phi_{\rm TV}<\phi<\phi_{\rm FV}$, slightly offset towards $\phi_{\rm FV}$.
At intermediate temperatures, the thermal contributions from self-interactions become Boltzmann-suppressed and can be ignored, and $V_{\rm eff}$ develops a second minimum at smaller field values. Notably, due to $V_{\cal P}$, the large-field minimum approaching $\phi_{\rm FV}$ is the global minimum of $V_{\rm eff}$ at these temperatures, and the dilaton is attracted to the false vacuum.
@@
We require that the transition percolates (at plasma temperature $T_{\star}$) which, for a PT with constant $\Gamma$, requires $\alpha\lesssim 20$ [34], where $\alpha=\Delta V/\rho_{\rm rad}(T_{\rm n})$. For such transitions, the inverse duration of the PT, $\beta$, is set by $\alpha$ and satisfies $38\gtrsim\beta/H_{*}>$ 3 [34]. As the energy density $\Delta V$ is channelled into the Standard Model, the plasma is reheated to the temperature $T_{\rm rh}=(1+\alpha)^{1/4}T_{n}$, here assuming $T_{\star}\approx T_{n}$. We restrict our discussion to $T_{\rm rh}\lesssim T_{\rm QCD}^{\rm TV}$, so that the hadrons do not deconfine (and then undergo a ‘second’ QCD transition, this time a cross-over).
The subsequent cosmology and observational predictions of this scenario depend on the speed at which the bubbles expand. As we show in the Supplementary Material [35], the bubble walls reach ultrarelativistic terminal velocities before bubbles collide.
Gravitational wave signal — As the relativistic dilaton bubbles grow large, collide, and percolate, they produce a stochastic gravitational wave (GW) background, potentially observable today [36, 37]. This is particularly interesting given the recent observation of the Hellings-Down correlation in Pulsar Timing Array (PTA) datasets [4, 5, 6, 7], providing evidence for a stochastic GW background in the nanohertz range. Such a background could potentially be explained astrophysically through supermassive black hole binary mergers [38], but this explanation has difficulties with accommodating the observed amplitude and spectrum [39, 40]. By contrast, several studies have found that GWs generated by a first-order PT provides a better fit to the data [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55] (see [56] for an alternative gravitational wave explanation generated by turbulent sources regardless of the order of the QCD PT).
Since the bubbles are expected to reach a terminal velocity before they collide, the main contribution to the GW signal in this scenario is from the dynamics of sound waves generated in the plasma [57, 58, 59, 60, 61] (the contribution to the signal coming from bubble wall collisions, which is relevant for runaway PTs, is discussed in the Supplementary Material [35]).
Estimates of the amplitude and spectrum of the GW background can be parametrised through rather simple formulas, depending on $T_{\star}\approx T_{n},~{}\alpha,~{}\beta$, and $v_{w}$ (cf. [62] and, for a recent review, [63]). The peak frequency is given by
$f_{\rm SW}=\frac{17.53~{}\text{nHz}}{v_{w}}\frac{\beta}{H_{*}}\frac{T_{n}}{100~{}\text{MeV}}\left(\frac{g_{*}}{61.75}\right)^{1/6}\;.$ (9)
@@
The corresponding GW energy spectra, $h^{2}\Omega_{\rm GW}$, give the fractional energy density per logarithmic frequency interval, rescaled by the squared Hubble parameter. The GW spectrum from sound waves is then
@@
Acknowledgments—We thank Mark Hindmarsh and Filippo Sala for stimulating discussions. This work was supported by the Swedish Research Council (VR) under grants 2018-03641, 2019-02337 and 2024-04289. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology).
Supplemental Material
Supplemental Material for ”Gravitational waves from a dilaton-induced, first-order QCD phase transition”
Appendix S1 Bubble wall velocity
In this section, we calculate the pressure exerted on the wall by the quark gluon plasma and determine the terminal speed of the bubble walls. Recall that we consider percolating phase transitions with time-independent $\Gamma$. The $\beta$-parameter is given by $\beta=\dot{I}(t_{\star})$, with $I(t)$ the average number of bubbles in the past comoving light cone [34], i.e.
[…]
where $a(t)$ denotes the scale factor, $r_{\rm com}(t_{\star},t)$ the comoving bubble radius and $t_{\star}$ the percolation time, defined by $I(t_{\star})=1$. As mentioned in the main text, for transitions that percolate, $0<\alpha\lesssim 20$, which corresponds to $38\gtrsim\beta/H_{*}>$ 3
The bubble wall velocity is determined by the balance of the driving pressure, $\mathcal{P}_{\rm dr}\equiv\Delta V$, and the pressure caused by friction from the surrounding plasma, $\mathcal{P}_{\rm fr}$. The calculation of bubble wall velocities in confining, first-order PTs is an active area of research [75, 76, 77, 78, 79, 79, 80, 81, 82, 83, 84, 85, 86, 87]. In this work, the driving pressure is generated by the dilaton potential instead of the confinement dynamics. The expanding bubbles are in the ballistic limit, meaning that the particles in the plasma do not interact significantly over the width of the wall, which in the plasma frame is given by $L_{w}/\gamma_{w}\sim 1/(m_{\phi}\gamma_{w})$ where $\gamma_{w}$ the Lorentz factor of the wall. The ballistic limit holds when $L_{w}/\gamma_{w}\ll L_{\rm MFP}$, where $L_{\rm MFP}$ is the mean-free path of the particles in the quark-gluon plasma outside the bubbles [88]. Parametrically, we expect $L^{-1}_{\rm MFP}\sim\alpha^{2}_{\rm s}(\phi_{\rm FV})\;T_{n}$ from $2\rightarrow 2$ scatterings. Since QCD remains perturbative in the false vacuum, the ballistic approximation holds for all values of $\gamma_{w}$.
$\mathcal{P}_{1\to 1}=\frac{T_{n}^{2}}{48}\left(g_{q}^{*}\Delta m_{q}^{2}+2g_{g}^{*}\Delta m_{g}^{2}\right)<\alpha\rho_{\rm rad}=\mathcal{P}_{\rm dr}\;,$ (15)
where $g_{q}^{*},\,g^{*}_{g}$ denote, respectively, the effective degrees of freedom for quarks and gluons in the plasma. The inequality assumes $\alpha\gtrsim 3\cdot 10^{-3}\,(\Lambda_{\rm QCD}/T_{n})^{2}$, which holds across the most relevant parameter space. Thus, one-to-one processes alone do not suffice to stop the runaway behaviour [89] 444 A more accurate description of friction in a confinement phase transition is given in [90], and results in the scaling $\mathcal{P}_{\rm friction}=\mathcal{P}_{\rm BS}\sim\sum_{i\in BS}\frac{m_{i}^{2}T_{n}^{2}}{24}$, which coincides with that of [89], used here. .
Gluon radiation is an additional source of friction. Here, we model the pressure from gluon emission at the wall interface by the equivalent expression for $W$ and $Z$ boson emission when entering bubbles of the low-temperature phase in models where the electroweak PT is first order [91] (we explore an alternative scenario with runaway walls in the next section). Adopting the results of [91, 62], the gluon emission pressure is given by
$\mathcal{P}_{1\rightarrow 2}\simeq N_{g}\Delta m_{g}\gamma\;\alpha_{s}(\phi_{w})T_{n}^{3}\;,$ (16)
with $\phi_{w}$ a value of the dilaton on the wall and $N_{g}=8$ is the number of gluons. Here, we have assumed that QCD remains perturbative at the wall, i.e. for an intermediate value of $\phi$. This is typically justified since $\Lambda_{\rm QCD}(\phi_{w})\ll\Lambda^{\rm TV}_{\rm QCD}$. The pressure contribution $\mathcal{P}_{1\rightarrow 2}$ grows with $\gamma$ and eventually catches up with $\Delta V$, leading to an ultra-relativistic terminal velocity with gamma factor  [^5] Another potential friction source is bound state formation from string fragmentation [77]. However, due to insignificant supercooling, we expect confinement to happen among quark pairs rather than quarks forming flux-tubes attached to the bubble wall, and string fragmentation to be unimportant. Even in this scenario though, the leading and next-to-leading order friction from gluons has similar dependence on $\gamma$
By comparing the size of the bubbles when reaching the terminal velocity to their separation $\beta^{-1}$, it is clear that the terminal velocity is reached well before bubbles collide.
Appendix S2 Gravitational waves from runaway walls
@@
In the main text, we have discussed the GW spectra in the well-motivated scenario where the dilaton bubble walls reach a terminal velocity before colliding. It is instructive to explore the alternate scenario in which the bubble walls ‘run away’. Such a scenario can be realised in models that yield sufficiently small QCD coupling at the bubble wall interface along with large values of $\beta/H_{*}$ so that bubbles never reach their terminal velocity before collision. When this happens, the dominant contribution to the gravitational wave spectrum is given by bubble wall collisions. As discussed in the main text, quantum phase transitions with $\Gamma\simeq$ constant yield $\beta/H_{*}\lesssim 8$, which is not very large. However, if the dilaton phase transition is for some reason driven by thermal fluctuations, larger values of $\beta/H_{*}$ can be attained, triggering runaway behaviour.
@@
Figure S1: Gravitational wave spectra from a dilaton-induced QCD PT overlaid with NANOGrav 15-year free-spectrum constraints (gray violins), including the contribution from bubble wall collisions. Solid lines show the spectra in the regime where the energy is carried by bubble walls, while dashed lines correspond to the regime where the energy is converted into sound waves, for comparison. Left: Spectra for transitions at fixed temperature $100\,\rm MeV$ with varying strength, $\alpha=0.7$, $3$, $6.8$. Right: Spectra for fixed strength $\alpha=6.8$ and varying temperature, $T_{n}=50$, $75$, $100~{}{\rm MeV}$.
@@
$h^{2}\Omega_{\rm coll.}(f)=1.96\times 10^{-5}\left(\frac{H_{*}}{\beta}\right)^{2}\left(\frac{\kappa_{\rm coll.}\alpha}{1+\alpha}\right)^{2}\left(\frac{61.75}{g_{*}(T_{\star})}\right)^{1/3}\left(\frac{0.11v_{w}^{3}}{0.42+v_{w}^{2}}\right)\left(\frac{3.8\left(f/f_{\rm coll.}\right)^{2.8}}{1+2.8\left(f/f_{\rm coll.}\right)^{3.8}}\right)\;,$ (18)
with peak frequency given by [59]
$f_{\rm coll}=\left(\frac{9.44~{}\text{nHz}}{1.8-0.1v_{w}+v_{w}^{2}}\right)\frac{\beta}{H_{*}}\frac{T_{n}}{100~{}\text{MeV}}\left(\frac{g_{*}}{61.75}\right)^{1/6}\;.$ (19)
Here $v_{w}=1$ and $\kappa_{\rm coll}$ is energy fraction that is converted into bubble walls. The corresponding spectra are shown in Fig. S1, where we use several benchmark values of the parameters. In the right panel we vary the transition strength ($\alpha=0.7$, $3$ and $6$) and set $T_{n}=100~{}{\rm MeV}$, while in the left panel the temperature is varied ($T_{n}=50~{}{\rm MeV}$, $75~{}{\rm MeV}$ and $100~{}{\rm MeV}$) and $\alpha=6.8$. The spectra are overlaid with the posterior distributions from the NANOGrav 15-year free-spectrum analysis [4], shown as gray violins. For solid lines we assume that all of the energy is in the form of colliding bubble walls, i.e. $\kappa_{\rm coll}=1$. The dashed lines in Fig. S1 show the corresponding spectra arising from sound waves, which were discussed in the main text.